Please ignore the circled choices. I am double checking my work.

Using the leters in -BANANA, since there are 3 As, and 2 Ns, the mumber of Set o nt arrangements of the letters is: a) 30 12. 60 e 120 720 one of these 13. You desidl to wach 4 hows if heo ten en you select the 4 shows you will watch? 32 0) 256 d) 1680 ) mone of these 14. Find the mean of a binoenial distribution where n-4 and p- 5 a) 0 b) 1 c) 1.5 2 ,04 15. Find the standard deviation of a binomial distribution where n-4 and p-s a) o c) 1.5 d 2 e) 4 16. Given a binomial distribution with 11 events and the peobability of success for an event oqal to 325 whats the probability of getting at least 4 saccesses? д) .2867 b)A552 _ c) .5448 d) .8029 ()none of these
Answer #1

12
different 6 letters words with same letters as BANANA = 6! / (3!*2!) = 720 / (6*2) = 60 (option B)

13
4 shows out of 8 shows can be chosen in 8C4 or 8! / (4!)(8-4)! = 70 (option B)

14
mean of a binomial distribution = n*p = 4*(0.5) = 2 (option C)

15
standard deviation of a binomial distribution = npq = 4*0.5*0.5 = 1 {where q = 1-p}

16 n = 11, p = 0.25, P(X>= 4)=0.28669548034 (option A)

please upvote if this helped